A company issues $15600000, 9.8%, 20-year bonds to yield 10% on January 1, 2017. Interest is paid on June 30 and December 31. The proceeds from the bonds are $15332318. Using effective-interest amortization, what will the carrying value of the bonds be on the December 31, 2017 balance sheet?
Answer #1

Answer : Bond Carrying Value = 15,345,702

Discount on Bonds Payable = 15600000-15332318 = 267,682

Semiannual Bond Discount Amortization = 267,682/40 =6692

Bond Carrying Value on December 31,2017 = 15332318+(6692*2) = 15,345,702 (Answer)